a bus is moving towards a huge wall with a velocity 5m/s. the driver soudes a horn of frequency 200 Hz. the frequency of beat heard by a passenger of the bus will be( speef of sound in air 342m/s)

Grade:12

2 Answers

Vikas TU
6 years ago
In the event that the sound achieves the spectator in the wake of being reflected from a stationary surface and the medium is additionally stationary, the picture of the source in the reflecting surface will turn into the wellspring of the reflected sound. 
v = v [c – v0/c – vs] 
v0, versus are +ve on the off chance that they are guided from source to the eyewitness and – ve in the event that they guided from spectator to source. 
v = 200 [342 – (- 5)/342 – 5] = 337 m/s 
∴ Beat recurrence = 337 – 200 = 137 Hz
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

In the event that the sound achieves the spectator in the wake of being reflected from a stationary surface and the medium is additionally stationary, the picture of the source in the reflecting surface will turn into the wellspring of the reflected sound.
v = v [c – v0/c – vs]
v0, versus are +ve on the off chance that they are guided from source to the eyewitness and – ve in the event that they guided from spectator to source.
v = 200 [342 – (- 5)/342 – 5] = 337 m/s
∴ Beat recurrence = 337 – 200 = 137 Hz

Thanks and Regards